# Transferable Tactile Transformers for Representation Learning Across Diverse Sensors and Tasks

### v2

@@ 3 The FoTa Dataset
Figure 2: FoTa dataset visualizations. (a) We show the mixture and distribution of constituent datasets, sensors, and tasks of the FoTa dataset. Note that not all tasks are utilized in the training of T3. (b) We visualize one tactile image from each constituent sensor in FoTa.
@@ 4 Heterogeneous Tactile Learning with T3
The training process of T3 is split into two pre-training phases and one optional fine-tuning phase. Pre-training is divided into two phases to improve data utilization and to attain both local fine-grained understanding and global semantic understanding. Phase 1 focuses on more pixel-level understanding using self-supervised learning, and it is able to utilize all data from the FoTa dataset. Phase 2 focuses on semantic understanding and it requires suitable task labels that only a subset of the FoTa dataset processes.
@@ 5 Experiments and Discussions
In this section we discuss the evaluation performance of T3 with numerous ablations and benchmarks. For both pre-training I and II, we use all data from the FoTa dataset except the additional ones collected on the gantry system in-house, which are saved for fine-tuning and performance benchmarking. We aim to answer the following questions in this section:
@@
Figure 3: Experiments on Task Performance and Transferability of T3. (a) Eval performance with different network sizes and different training schemes. (b) Eval performance with different masking ratios during Pre-training I. (c) Transferability test on a classification task. (d) Transferability test on a pose estimation task.
@@ 5.1 The worthiness of pre-training
The fine-tuning and evaluation task for this section is an object classification task (6 categories) for 2 sensors (GelSight Wedge, GelSight Mini). Averaged top-1 classification accuracy from the two sensors is reported.
@@
We constructed 4 network variations, tiny (12M params), small (45M params), medium (174M params), and large (308M params). For each network configuration, we tested 3 training schemes: train from scratch, fine-tune from pre-training I, and fine-tune from pre-training I & II. The evaluation accuracy of the classification task is shown in Fig. 3.a. We observed that pre-training significantly improved the evaluation performance with all network configurations with an average improvement of 31%. We also observed that the evaluation performance improved with larger networks, where large’s classification accuracy was 19% higher than that of tiny. However, the performance difference between base and large was insignificant.
@@ 5.2 Zero-shot transferability of pre-trained T3 to new sensors or tasks
We tested to what extent a pre-trained T3 can be transferred to a new sensor or downstream task with and without fine-tuning. We picked two sensors (GelSight Svelte, DensetTact2.0), two tasks (object classification, pose estimation between two overlapped tactile images). No combinations of the two sensors and the two tasks have been jointly trained during pre-training, i.e. both sensors have only been trained for different tasks, and both tasks have only been trained with other sensors. Results on the 4 combinations are shown in Fig. 3.c and Fig. 3.d. Zero-shot transfer yielded only minor improvements over random guesses in the object classification task; however, it demonstrated significant improvement over dataset averages in the pose estimation task. We further fine-tuned the networks with a small amount of 2,000 data points. The classification accuracy immediately bounced up to 17% with fine-tuning, and the pose estimation root mean squared error (RMSE) was reduced by 5.5mm for DenseTact2.0. On the GelSight Svelte sensor, the pose estimation errors before and after fine-tuning were nearly identical and close to optimal. These results indicate that zero-transferability can be achieved in certain sensor-task pairings, and that better outcomes can be attained with minimal fine-tuning.
@@ 5.3 T3 in long-horizon manipulation tasks
Our experiment setup consists of two GelSight Wedge tactile sensors mounted on a parallel jaw gripper, which is attached to the end-effector of a 7-DoF Franka Emika Panda robot, a PCB board fixed at a known location on the workbench, the 3 electronics parts, and a RGB camera on the side of the workbench. An illustration of the setup is shown in Fig. 4 a and the parts are shown in Fig. 4.b. Data is collected by controlling the robot in "guide mode", i.e. under zero stiffness control, and a human operator manually grabs and moves the end-effector until the part is fully inserted. Randomization is applied at the grasping pose during data collection.
@@ 6 Limitations and Future Works
One limitation of the current training pipeline is that the FoTa dataset is unbalanced. Data collected with the 2 most popular sensors constitutes over 50% of the entire dataset. Therefore, the trained policy could also be biased towards those more popular sensors. Another limitation is that the current architecture of T3 focuses on per-image encoding and T3 is trained with explicit labels during pre-training II and fine-tuning. Representation learning on tactile image sequences with sparse or implicit labels can be a future direction of this work.
@@ Appendix A Appendix
A.1 Public datasets aggregated in the FoTa dataset
@@ A.2 Data format
A.3 Data collection platforms
A.3.1 7-DoF robotic platform
@@ A.4 Network configurations
The decoder architecture varies based on the tasks.
  • The reconstruction decoder for masked auto encoding during pre-training I is also a ViT, with a MLP ratio of $4.0$, an embedding dimension of $512$, $16$ heads, and $8$ layers. No pooling is performed after the trunk for this decoder.
@@
Table 2: Network Sizes
$D_{enc}$ $H_{enc}$ $L_{enc}$ $L_{tru}$ Params incl. all encoders and decoders
T3-tiny 192 3 3 9 12M
T3-small 384 6 3 9 45M
T3-medium 768 12 3 9 174M
T3-large 1024 8 3 9 308M

### v3

@@ 3 The FoTa Dataset
Figure 2: FoTa dataset visualizations. (a) We show the mixture and distribution of constituent datasets, sensors, and tasks of the FoTa dataset. Note that not all tasks are utilized in the training of T3. (b) We visualize one tactile image from each constituent sensor in FoTa. Note that in the training of T3, similar sensors share encoders. For example, GelSight17 var. {1-4} share the same encoder, and GelSight var. {1-2} share the same encoder.
@@ 4 Heterogeneous Tactile Learning with T3
The training process of T3 is split into two pre-training phases and one optional fine-tuning phase. Pre-training is divided into two phases to improve data utilization and to attain both local fine-grained understanding and global semantic understanding. Phase 1 focuses on more pixel-level understanding using self-supervised learning, and it is able to utilize all data from the FoTa dataset. Phase 2 focuses on semantic understanding and it requires suitable task labels that only a subset of the FoTa dataset possesses.
@@ 5 Experiments and Discussions
In this section we discuss the evaluation performance of T3 with numerous ablations and benchmarks. For both pre-training I and II, we use all data from the FoTa dataset except data collected on the gantry system in-house, including the object classification task and the 3D pose estimation task. Those data are therefore not included for pre-training the foundation model and are set aside for evaluation and fine-tuning only. We aim to answer the following questions in this section:
@@
Figure 3: Experiments on Task Performance and Transferability of T3. (a) Eval performance with 4 network sizes (tiny, small, medium, large), 3 training schemes (train from scratch, fine-tune from pre-train 1, fine-tune from pre-train 2), and 2 amounts of fine-tuning data (half data and full data) (b) Eval performance with different masking ratios during Pre-training I. (c) Transferability test on a classification task. (d) Transferability test on a pose estimation task.
@@ 5.1 The worthiness of pre-training
The fine-tuning and evaluation task for this section is an object classification task (6 categories) for 2 sensors (GelSight Wedge, GelSight Mini) with around 3,300 data points for each sensor. Averaged top-1 classification accuracy from the two sensors is reported.
@@
We conducted the classification experiment with all combinations of the following: (a) 4 network variations: tiny-12M, small-45M, medium-174M, and large-308M; (b) 3 training schemes: train from scratch, fine-tune from pre-training I, and fine-tune from pre-training I & II; (c) 2 amounts of training data: half data (approx. 1,650 training data) and full data (approx. 3,300 training data). All 24 experiments were tested on the same validation dataset, which was separated from the training set. The evaluation accuracy is shown in Fig. 3.a. We observed that: (a) Pre-training significantly improved the evaluation performance with all network configurations with an median improvement of 24%. (b) The evaluation performance improved with larger networks, where large’s classification accuracy was 19% higher than that of tiny. However, the performance difference between medium and large was insignificant. (c) When fine-tuned with half of the fine-tuning data, pre-trained networks performed better than trained from scratch networks, and larger pre-trained models performed better than smaller pre-trained models. In fact, for medium and large, the performances between fine-tuned with half data and fine-tuned with full data were very close, signaling those models generalized better thus they require less data to fine-tune to a novel task.
More qualitatively analysis, including visualizations of the encoder attention maps and the trunk attention maps, can be found in Appendix. A.1.
@@ 5.2 Zero-shot transferability of pre-trained T3 to new sensors or tasks
We tested to what extent a pre-trained T3 can be transferred to a new sensor or downstream task with and without fine-tuning. We picked two novel sensors (GelSight Svelte, DensetTact2.0) and two seen tasks (object classification, pose estimation between two overlapped tactile images). The two seen tasks were pre-trained with data from GelSight Wedge, GelSight Finray, and GelSight Mini. During this experiment, we use the encoder pre-trained for GelSight Wedge as the encoder for GelSight Svelte, and GelSight Mini’s encoder in place of DenseTact2.0’s encoder. Results on the 4 combinations are shown in Fig. 3.c and Fig. 3.d. Zero-shot transfer yielded only minor improvements over random guesses in the object classification task; however, it demonstrated significant improvement over dataset averages in the pose estimation task. We further fine-tuned the networks with a small amount of 2,000 data points. The classification accuracy immediately bounced up to 17% with fine-tuning, and the pose estimation root mean squared error (RMSE) was reduced by 5.5mm for DenseTact2.0. On the GelSight Svelte sensor, the pose estimation errors before and after fine-tuning were nearly identical and close to optimal. These results indicate that zero-transferability can be achieved in certain sensor-task pairings, and that better outcomes can be attained with minimal fine-tuning.
@@ 5.3 T3 in long-horizon manipulation tasks
Our experiment setup consists of two GelSight Wedge tactile sensors mounted on a parallel jaw gripper, which is attached to the end-effector of a 7-DoF Franka Emika Panda robot, a PCB board fixed at a known location on the workbench, the 3 electronics parts, and a RGB camera on the side of the workbench. An illustration of the setup is shown in Fig. 4 a and the parts are shown in Fig. 4.b. Data is collected by controlling the robot in "guide mode", i.e. under zero stiffness control, and a human operator manually grabs and moves the end-effector until the part is fully inserted. At both data collection and inference time, the sockets are fixed, and we add randomizations to the part-in-hand pose and gripper starting pose before each insertion, both generated from a 3D zero-mean, $1mm$ standard deviation normal distribution.
@@ 6 Limitations and Future Works
One limitation of the current training pipeline is that the FoTa dataset is unbalanced. Data collected with the 2 most popular sensors constitutes over 50% of the entire dataset. Therefore, the trained policy could also be biased towards those more popular sensors. Another limitation is that the current architecture of T3 focuses on per-image encoding and T3 is trained with explicit labels during pre-training II and fine-tuning. Representation learning on tactile image sequences with sparse or implicit labels can be a future direction of this work. Furthermore, T3 is currently limited to learning representations for camera-based tactile sensors. Extending the FoTa dataset and the T3 framwork to non-camera-based sensors is another future line of work.
@@ Appendix A Appendix
A.1 Visualization of attention maps
Figure 5: Visualization of self-attention weights of the encoders and the trunk of pre-trained T3.
Top: the original tactile images. Middle: encoder attention weights applied on the original images as masks. Bottom: trunk attention weights applied on the original images as masks
We calculate the attention weights $attn_{i}=softmax(\frac{Q\cdot K}{scale})$ at layer $i$ of each self-attention layers for all encoders and the trunk. We then obtain a joint attention map, calculated as the product of the attention weights of all layers within each encoder or trunk, $map=\prod_{i}attn_{i}$. This joint attention map is then normalized and applied on the original tactile images as masks. A few tactile images as well as them applied with encoder masks and trunk masks are visualized in Fig. 5. Qualitatively, we observe that the encoder self-attentions highlight the region of the contact object, whereas the trunk self-attentions are more spread out and they cover other features like the markers on a sensor. This result seems to show that the encoder is more focused on the sensor agnostic regions, i.e. the "hot" regions of the contact, while the trunk also attends to sensor specific areas.
A.2 Public datasets aggregated in the FoTa dataset
@@ A.3 Data format
Table 1: Network Sizes
$D_{enc}$ $H_{enc}$ $L_{enc}$ $L_{tru}$ Params incl. all encoders and decoders
T3-tiny 192 3 3 9 12M
T3-small 384 6 3 9 45M
T3-medium 768 12 3 9 174M
T3-large 1024 8 3 9 308M
@@
A.4 Data collection platforms
A.4.1 7-DoF robotic platform
@@ A.5 Network configurations
The decoder architecture varies based on the tasks. The reconstruction decoder for masked auto encoding during pre-training I is also a ViT, with a MLP ratio of $4.0$, an embedding dimension of $512$, $16$ heads, and $8$ layers. No pooling is performed after the trunk for this decoder.